Hello, this is a question from my Algebra 2 textbook:
Find all values of c for which the equation has (a) two real solutions, (b) one real solution, and (c) two imaginary solutions.
x^2 - 2x + c = 0
Thanks!

Answer by jim_thompson5910(35256)

In the case of x^2 - 2x + c = 0, it is of the form ax^2 + bx + c = 0 where a = 1, b = -2 and c is unknown.


The equation ax^2 + bx + c = 0 has at most 2 solutions. The type and number of unique solutions are determined by the equation below

D = b^2 - 4ac

which is the discriminant formula. The rules are this

* If D > 0, then you'll have 2 real solutions that are distinct.

* If D = 0, then you'll have exactly 1 real solution (that's also rational).

* If D < 0, then you'll have 2 imaginary solutions (aka, complex solutions)


---------------------------------------------------------------------------

So from the rules above, we will have 2 real solutions when D > 0.

D > 0

b^2 - 4ac > 0 ... replace D with b^2 - 4ac (since D = b^2 - 4ac)

(-2)^2 - 4(1)c > 0 ... plug in a = 1, b = -2, and leave c alone

4 - 4c > 0

4 - 4c + 4c > 0 + 4c

4 > 4c

4c < 4

4c/4 < 4/4

c < 1

So if c < 1, then D > 0 which will lead to 2 distinct real solutions.

-------------------------------------------------------

Now let's move onto the second case: When D = 0


D = 0

b^2 - 4ac = 0

(-2)^2 - 4(1)c = 0

4 - 4c = 0

4 = 4c

4c = 4

c = 1

If c = 1, then you'll get exactly one real solution.

-------------------------------------------------------

You can probably guess what the last answer must be, but let's work through it

D < 0

b^2 - 4ac < 0

(-2)^2 - 4(1)c < 0

4 - 4c < 0

4 < 4c

4c > 4

c > 1

So if c > 1, then you'll get 2 imaginary solutions.